40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed May 25 16:56:44 2022, Paco, Configuration, BHD, IFO recovery - IMC alignment 
    Reply  Wed May 25 17:34:47 2022, yuta, Configuration, BHD, IFO recovery - IFO alignment Screenshot_2022-05-25_17-47-57.png
       Reply  Thu May 26 19:55:43 2022, yuta, Configuration, BHD, Oplevs centered, BHD DCPDs are now online elog_1Y2.JPGelog_BHD.JPGelog_box.JPGScreenshot_2022-05-26_17-37-27_IFOaligned_OplevCentered.pngScreenshot_2022-05-26_20-35-02.png
          Reply  Fri May 27 17:45:53 2022, yuta, Configuration, BHD, BHD camera installed, GRY aligned 
             Reply  Thu Jun 2 20:05:37 2022, yuta, Configuration, PSL, IMC input power recovered to 1W, some alignment works Before.JPGAfter.JPG
                Reply  Fri Jun 3 12:13:58 2022, Paco, Configuration, CDS, Fix RFM channels  SoGreen.png
                   Reply  Fri Jun 3 15:22:51 2022, yuta, Update, LSC, Both arms locked with POY/POX, IR beam centered on TMs with ASS IRBeamsOnTMs.JPGScreenshot_2022-06-03_15-03-51.png
                      Reply  Fri Jun 3 17:42:50 2022, yuta, Update, LSC, MICH locks with AS55_Q Screenshot_2022-06-03_17-41-55.png
                         Reply  Sun Jun 5 19:46:40 2022, Paco, Update, LSC, Fixed IMC Trans sum issue 
                         Reply  Mon Jun 6 13:35:11 2022, Paco, Update, LSC, First calibrated spectra of MICH at AS55 Q Screenshot_2022-06-06_17-30-16_MICHOLTF.pngCalibrated_MICH_ERR.pdf
                            Reply  Fri Jun 17 16:22:21 2022, yuta, Update, LSC, Actuator calibration of BS. ITMX, ITMY, updated MICH displacement spectra from c1cal 6x
                               Reply  Thu Jul 7 18:18:19 2022, yuta, Update, LSC, Actuator calibration of ETMX and ETMX Screenshot_2022-07-05_14-52-01_OLTF.pngScreenshot_2022-07-05_14-54-03_TF.pngScreenshot_2022-07-05_14-56-41_Ratio.png
                                  Reply  Thu Jul 7 18:22:12 2022, yuta, Update, LSC, Actuator calibration of MC2 using Yarm TF.pngMC2.png
                                     Reply  Fri Jul 8 16:18:35 2022, rana, Update, LSC, Actuator calibration of MC2 using Yarm 
                                     Reply  Mon Jul 18 17:07:12 2022, yuta, Update, LSC, x4.12 added to ETMX coil outputs to balance with ETMY 
          Reply  Thu Jun 16 18:23:15 2022, Paco, Configuration, BHD, Recovering LO beam in BHD DCPDs Screenshot_2022-06-16_18-29-14_BHDLObeamISBACK.png
          Reply  Tue Jun 21 14:17:50 2022, yuta, Configuration, BHD, BHD DCPDs re-routed to c1sus2 C1X07ADC1.JPGBHDDCPDs.JPG
             Reply  Tue Jul 19 16:00:34 2022, yuta, Configuration, BHD, Fast channels for BHD DCPDs now available in c1lsc but not in c1hpc Screenshot_2022-07-19_14-26-39_c1hpc.pngScreenshot_2022-07-19_14-24-49_c1lsc.pngScreenshot_2022-07-19_15-51-25_GreenGreen.png
                Reply  Thu Jul 21 21:50:47 2022, Tega, Configuration, BHD, c1sus2 IPC update 
                   Reply  Fri Jul 22 15:05:26 2022, Tega, Configuration, BHD, c1sus2 shared memory and ADC fix Screenshot_2022-07-22_15-04-33_DCPD.pngScreenshot_2022-07-22_15-12-19_models.pngScreenshot_2022-07-22_15-15-11_ERR.pngScreenshot_2022-07-22_15-32-19_GDS.png
                      Reply  Fri Jul 22 17:46:10 2022, yuta, Configuration, BHD, c1sus2 watchdog update and DCPD ERR channels Screenshot_2022-07-22_17-48-25.png
                         Reply  Mon Jul 25 17:58:10 2022, Tega, Configuration, BHD, c1sus2 IPC dolphin issue update Screen_Shot_2022-07-25_at_5.43.28_PM.pngScreen_Shot_2022-07-25_at_5.21.10_PM.pngScreen_Shot_2022-07-25_at_5.30.58_PM.pngScreen_Shot_2022-07-25_at_5.35.19_PM.png
                            Reply  Mon Aug 1 18:42:39 2022, Tega, Configuration, BHD, c1sus2 IPC dolphin issue update c1sus2_dolphin.pngfb1_dxamin_status.pngdolphin_num_mem_init2.png
                               Reply  Tue Aug 2 17:25:18 2022, Tega, Configuration, BHD, c1sus2 dolphin IPC issue solved c1lsc_IPC_status.pngFE_lsmod_dependencies_c1sus2_b4_after_iop_unpdate.png
                      Reply  Mon Jul 25 18:09:41 2022, Tega, Configuration, BHD, BHD Homodyne Phase control MEDM screen Screen_Shot_2022-07-25_at_6.12.08_PM.pngScreen_Shot_2022-07-25_at_6.18.09_PM.png
Message ID: 16888     Entry time: Fri Jun 3 15:22:51 2022     In reply to: 16887     Reply to this: 16889
Author: yuta 
Type: Update 
Category: LSC 
Subject: Both arms locked with POY/POX, IR beam centered on TMs with ASS 

[JC, Paco, Yuta]

We locked both Y and X arms with POY11 and POX11.
RFM fix (40m/16887) enabled us to use triggering using C1:LSC-TRY/X_OUT.
IR beam is now centered on TMs using ASS (for Yarm, ASS loops cannot be closed fully, so did it manually).

What we did:
 - Aligned both arms so that the beams are roughly centered at TMs using cameras.
 - Yarm lock was easy, but Xarm lock required gain tuning. Somehow, Xarm required x3 higher gain as follows, although the amplitude of POX11_I_ERR seems to be almost the same as POY11_I_ERR. I suspect it is something to do with power normalization matrix (TRX flashing is almost a double of TRY flashing).

C1:LSC-YARM_GAIN = 0.01
C1:LSC-XARM_GAIN = 0.03

 - Run ASS for Yarm. ASS loops cannot be closed fully using default feedback parameters. I guess this is because ITMY ULCOIL is not working (40m/16873). ASS demodulated signals were manually zero-ed by manually aligning ETMY, ITMY and PR3 (and some TT1 and TT2), except for demodulated signals related to ITMY. Beam on ITMY was centered just by using our eyes.
 - Run ASS for Xarm. It seemed to work well.
 - After this, TRX and TRY were as follows and beam positions on TMs were as attached.

C1:LSC-TRX_OUT ~0.95
C1:LSC-TRY_OUT ~ 0.58

(TRX is somehow lower than what we had yesterday... 40m/16886; TRX and TRY photodiode alignment was checked, but seems to be OK.)

 - Centered TMs and BS oplevs.

Next:
 - POX and POY demodulation phases are not fully optimized. Needs re-tuning.
 - Tweak GRX and GRY injection (restore GRY PZTs?)
 - Install ETMXT camera (if it is easy)
 - MICH locking
 - RTS model for BHD needs to be updated

Attachment 1: IRBeamsOnTMs.JPG  3.049 MB  Uploaded Fri Jun 3 16:27:46 2022  | Hide | Hide all
IRBeamsOnTMs.JPG
Attachment 2: Screenshot_2022-06-03_15-03-51.png  187 kB  Uploaded Fri Jun 3 16:28:14 2022  | Hide | Hide all
Screenshot_2022-06-03_15-03-51.png
ELOG V3.1.3-